OpenStudy (anonymous):

Eddie has saved up $45 to purchase a new camera from the local store. The sales tax in his county is 7% of the sticker price. Write an equation and solve it to determine the value of the highest priced camera Eddie can purchase with his $45, including the sales tax. Round your answer to the nearest penny. x - 0.07x = 45; $48.39 0.7x = 45; x = $64.29 7x = 45; x = $6.43 x + 0.07x = 45; x = $42.06

OpenStudy (anonymous):

@ganeshie8 @Hero @ParthKohli

OpenStudy (jdoe0001):

so if you go the store with $45 and say the camera has a sticker price of $45, can you afford it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well, no because it will include the tax

OpenStudy (jdoe0001):

well.. what about when you show up at the counter, and the clerk asks you for the tax =) right

OpenStudy (anonymous):

and if the base price is 45 but the tax is let's say 3.50? then i can't afford it so yeah

OpenStudy (jdoe0001):

so... you can't do with $45.... you really need THAT $45 to be, the STICKER PRICE of the camera PLUS TAX already

OpenStudy (anonymous):

ok so i need to multiply 45 by 0.07

OpenStudy (anonymous):

the tax will equal 3.15 right?

OpenStudy (jdoe0001):

so... whatever the camera may cost, plus tax, should be 45 so let's use see what 7% of something is so \(\large { {\color{red}{ 7}}\%\ of \ camera=\cfrac{{\color{red}{ 7}}}{100}\times camera \iff tax\qquad thus \\ \quad \\ \begin{array}{rllll} camera\\ +\frac{{\color{red}{ 7}}}{100}\times camera\\ \hline\\ total\ price \end{array} }\) and then we know that \(\bf \cfrac{7}{100}\to 0.07\qquad thus \qquad \cfrac{7}{100}camera\to 0.07camera\) so one can say that \(\large \begin{array}{rllll} camera\\ +0.07\times camera\\ \hline\\ 45 \end{array}\)

OpenStudy (jdoe0001):

and if we use "x" for the camera then one can say that the price of it will be "x" and the equation then 0.07x+x=45

OpenStudy (anonymous):

So it's D

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thanks! :)

OpenStudy (jdoe0001):

yw